# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ B.3 Stability Results
By (B.25), it holds that
$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}$ $=\left|\int_{{}^{d_{u}}}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\right|$
$\leqslant\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Now, since $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)/\mathsf{P}\mu(u)$ is a Gaussian density with covariance $\Gamma$, which is bounded uniformly from above by $\bigl{(}(2\pi)^{d_{y}}\det(\Gamma)\bigr{)}^{-1/2}$, we also have that
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$ $=\beta_{\mu}(y)+\int_{{}^{d_{u}}}|u|^{2}\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u$
$\leqslant\beta_{\mu}(y)+\int_{{}^{d_{u}}}\frac{|u|^{2}\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}\,\mathrm{d}u$
$=\beta_{\mu}(y)+\frac{\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}.$
We proceed by noting the result of Lemma B.1, which gives that the first and second moments of $\mathsf{P}\mu$ are bounded from above by a constant which depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\Sigma$.
@@
Suppose that Assumption H is satisfied. Then, for all $R\geqslant 1$, it holds that for any $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, there is $L_{\mathsf{T}}:=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$, such that
@@
By the results of Lemma B.2, it holds that the probability measure $\mathsf{Q}\mathsf{P}\mu$ belongs to $\mathcal{P}_{\widetilde{R}}(^{d_{u}}\times^{d_{y}})$ for some $\widetilde{R}\geqslant 1$. Define
@@
We denote by $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ the affine maps corresponding to evaluation of covariance information at the probability measures $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$, respectively. Namely, letting $K=\mathcal{C}(\pi)$, $S=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$, we write
@@
Before bounding each term on the right-hand side of B.26, we derive some inequalities that we will use in the proof. It holds that
$\lVert A_{\pi}\rVert\leqslant\lVert K_{uy}\rVert\lVert K_{yy}^{-1}\rVert\leqslant\lVert K\rVert\lVert K^{-1}\rVert\leqslant r^{4},$ (B.27)
and similarly for $A_{p}$. This can easily be seen by noting that the matrix norm (induced by the Euclidean vector norm) of any submatrix is bounded from above by the norm of the full matrix. Using this bound again and assuming without loss of generality that $r>1$, we deduce that
$\lVert A_{\pi}-A_{p}\rVert=\lVert K_{uy}K_{yy}^{-1}-S_{uy}S_{yy}^{-1}\rVert$ $\leqslant\lVert(K_{uy}-S_{uy})K_{yy}^{-1}\rVert+\lVert S_{uy}\left(K_{yy}^{-1}-S_{yy}^{-1}\right)\rVert$
$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}-S_{yy}^{-1}\rVert\Bigr{)}$
$=r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}(S_{yy}-K_{yy})S_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant r^{2}\Bigl{(}\lVert K_{uy}-S_{uy}\rVert+\lVert K_{yy}^{-1}\rVert\lVert S_{yy}-K_{yy}\rVert\lVert S_{yy}^{-1}\rVert\Bigr{)}$
$\leqslant 2r^{6}\lVert K-S\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.28)
where the last line follows from the result in Lemma B.6.
@@
Let $f\colon^{d_{u}}\to$ be a function such that $\lvert f\rvert\leqslant g$ with $g(u)=1+\lvert u\rvert^{2}$. It follows from the definition of the pushforward measure that
$\bigl{\lvert}\mathscr{T}^{\pi}_{\sharp}\pi[f]-\mathscr{T}^{\pi}_{\sharp}p[f]\bigr{\rvert}=\bigl{\lvert}\pi[f\circ\mathscr{T}^{\pi}]-p[f\circ\mathscr{T}^{\pi}]\bigr{\rvert}.$
For any $(u,y)\in^{d_{u}}\times^{d_{y}}$, we have that
$\bigl{\lvert}f\circ\mathscr{T}^{\pi}(u,y)\bigr{\rvert}$ $=\left\lvert f\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}\right\rvert\leqslant g\Bigl{(}u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\Bigr{)}$
$=1+\bigl{\lvert}u+A_{\pi}(y^{\dagger}-y)\bigr{\rvert}^{2}\leqslant 1+3\lvert u\rvert^{2}+3\lvert A_{\pi}y^{\dagger}\rvert^{2}+3\lvert A_{\pi}y\rvert^{2}$
$\leqslant 3\left(1+\lvert A_{\pi}y^{\dagger}\rvert^{2}\right)\max\left\{1,\lVert A_{\pi}\rVert^{2}\right\}g(u,y).$
Hence, by (B.27) we conclude that
@@
Let $f\colon^{d_{u}}\to$ again be a function such that the inequality $\lvert f\rvert\leqslant g$, with $g(u)=1+\lvert u\rvert^{2}$, is satisfied. We note that
@@
The right-hand side may be rewritten as
$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f\left(u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+A_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert.$
A change of variable yields
$\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\bigl{(}f(u+A_{\pi}z)-f(u+A_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f(v)\Bigl{(}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{)}\mathrm{d}v\,\mathrm{d}z.$ (B.30)
Since $\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty$ by assumption, from Lemma A.3 it follows that
@@
$\qquad\leqslant C\lvert A_{\pi}z-A_{p}z\rvert\cdot\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\cdot$
$\qquad\qquad\cdot\frac{1}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}.$
We apply this inequality to bound for fixed $z\in^{d_{y}}$ the inner integral in (B.30). Keeping only the terms that depend on $v$, we obtain that
$\int_{{}^{d_{u}}}\lvert f(v)\rvert\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\,\mathrm{d}v$
$\qquad\leqslant\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v+\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\mathrm{d}v.$
Since $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, proceeding by substitution gives that this expression is bounded from above by
@@
where the last inequality follows from (B.27). The remaining integral in the $z$ variable can be bounded in a similar way:
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma 28], where a complete proof is given. ∎
@@
In [le2009large] machinery is established to prove Monte Carlo error estimates between the finite particle ensemble Kalman filter and its mean field limit. We use such results as a component in proving Theorem 2.4 and, in so-doing, explicit dependence on moments must be tracked. In this section we detail proofs of the result in  [le2009large, Proposition 4.4] and [le2009large, Theorem 5.2] , following the analysis closely and, in addition, tracking dependence on moments. This leads to the following error estimate stating the desired Monte Carlo error estimate.
@@
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H5). We assume that the vector field $h$ is a linear transformation, i.e. that Assumption (V2) is satisfied. Furthermore, if the vector field $\Psi$ additionally satisfies $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, then for all $\phi$ satisfying Assumption P1, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu^{\rm EK}_{0}),J,R_{\phi},L_{\phi},\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\Sigma,\Gamma\bigr{)}$ where $q\coloneqq\max\{16\cdot 3^{J-1},8(\varsigma+1)3^{J-1}\}$ such that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{\rm EK}_{J}[\phi]\Bigr{|}^{2}\right)^{1/2}\leqslant\frac{C}{\sqrt{N}}.$
@@
To prove the proposition, we apply the coupling argument used in [le2009large]. Using similar notation to the one in [le2009large], to the interacting $N$-particle system $\{u^{(i)}_{j}\}_{n=1}^{N}$ evolving according to the ensemble Kalman dynamics (2.5) , we couple $N$ copies of the mean field dynamics $\{\overline{u}^{(j)}_{j}\}_{n=1}^{N}$ evolving according to the mean field ensemble Kalman dynamics (2.1). The mean field replicas are synchronously coupled to the interacting particle system, in the sense that they are initialized at the same initial condition and driven by the same noises; namely, the two particle systems are initialized at i.i.d. samples $u_{0}^{(i)}=\overline{u}_{0}^{(i)}\sim\mu_{0}^{\rm EK}$ for $i=1,\ldots,N$. For simplicity of notation, the forecast particles are denoted by the letter $v$, and we drop the hat notation from the forecast and simulated observations. Furthermore, we add a bar $\overline{\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}}$ to all the variables related to the synchronously coupled mean field particles, including the probability measures $\overline{\pi}^{\rm EK}_{j}$. We define for $\pi\in\mathcal{P}(^{d_{u}\times d_{y}})$ the Kalman gain
@@
Note that the synchronously coupled particles are independent and identically distributed. With this set-up we proceed with the proof. We first define
$D_{j}^{N,p}\coloneqq\Bigl{(}\mathbb{E}\left\lvert u^{(i)}_{j}-\overline{u}^{(i)}_{j}\right\rvert^{p}\Bigr{)}^{1/p}.$
We now let $\phi$ satisfy Assumption P1. By applying the triangle inequality, we deduce that

### v2

@@ B.3 Stability Results
In view of (B.23), it holds that
$\bigl{\lvert}\alpha_{\mu}(y)-\beta_{\mu}(y)\bigr{\rvert}\leqslant\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu).$
Now, since $\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})/\mathsf{P}\mu(u)$ defines a Gaussian density which has covariance $\Gamma$ and is bounded uniformly from above by $\bigl{(}(2\pi)^{d_{y}}\det(\Gamma)\bigr{)}^{-1/2}$, we also have that
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\mathsf{Q}\mathsf{P}\mu(u,y)\,\mathrm{d}u\leqslant\int_{{}^{d_{u}}}\frac{\left(1+|u|^{2}\right)\mathsf{P}\mu(u)}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}\,\mathrm{d}u=\frac{1+\operatorname*{tr}\bigl{(}\mathcal{C}(\mathsf{P}\mu)\bigr{)}+\left\lvert\mathcal{M}(\mathsf{P}\mu)\right\rvert^{2}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma})}.$
By the moment bounds in Lemma B.1, the right-hand side is bounded from above by a constant which depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\Sigma$.
@@
Suppose that Assumption H is satisfied. Then, for all $R\geqslant 1$, it holds that for any $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, there is $L_{\mathsf{T}}:=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\Sigma,\Gamma\bigr{)}$, such that
@@
By the results of Lemma B.2, it holds that $\mathsf{Q}\mathsf{P}\mu\in\mathcal{P}_{\widetilde{R}}(^{d_{u}}\times^{d_{y}})$ for some $\widetilde{R}(\mathcal{M}_{2}(\mu),\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma)\geqslant 1$. Using analogous notation to the one found in the proof of [carrillo2022ensemble, Lemma B.10] we define
@@
Letting $K=\mathcal{C}(\pi)$, $S=\mathcal{C}(p)$ and $y^{\dagger}=y^{\dagger}_{j+1}$, we also define the affine maps $\mathscr{T}^{\pi}$ and $\mathscr{T}^{p}$ corresponding to use of covariance information at the probability measures $\pi$ and $p=\mathsf{Q}\mathsf{P}\mu$,
@@
In the following steps we will separately bound the two terms on the right hand side of (B.24). Before proceeding we outline two auxiliary bounds that will be used in the rest of the proof. These bounds are identical to the ones found in the proof of [carrillo2022ensemble, Lemma B.10]; we include statements here for expository purposes. Noting that the operator 2-norm of any submatrix is bounded from above by the operator 2-norm of the full matrix, we observe (see [carrillo2022ensemble, Equation (B.23)]) that
$\lVert A_{\pi}\rVert\leqslant\lVert K_{uy}\rVert\lVert K_{yy}^{-1}\rVert\leqslant\lVert K\rVert\lVert K^{-1}\rVert\leqslant r^{4}.$ (B.25)
Note that the above bound similarly holds for $A_{p}$. Using (B.25) and assuming without loss of generality that $r>1$, we deduce (see [carrillo2022ensemble, Equation (B.24)]) that
$\lVert A_{\pi}-A_{p}\rVert\leqslant\lVert(K_{uy}-S_{uy})K_{yy}^{-1}\rVert+\lVert S_{uy}\left(K_{yy}^{-1}-S_{yy}^{-1}\right)\rVert\leqslant 2r^{6}\lVert K-S\rVert\leqslant 2r^{6}(1+2r)\,d_{g}(\pi,p),$ (B.26)
where the second inequality follows again from fact that the 2-norm of any submatrix is bounded from above by the 2-norm of the full matrix, while the last inequality follows from the result in Lemma B.6.
@@
With an identical argument to the one used in the proof of [carrillo2022ensemble, Lemma B.10] we have that
@@
Let $f$ again satisfy $\lvert f\rvert\leqslant g$. We note that
@@
The last term may be expressed as
$\Bigl{\lvert}p[f\circ\mathscr{T}^{\pi}-f\circ\mathscr{T}^{p}]\Bigr{\rvert}$ $=\left\lvert\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\Bigl{(}f\left(u+A_{\pi}\bigl{(}y^{\dagger}-y\bigr{)}\right)-f\left(u+A_{p}\bigl{(}y^{\dagger}-y\bigr{)}\right)\Bigr{)}\,p(u,y)\,\mathrm{d}u\,\mathrm{d}y\right\rvert$
$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\bigl{(}f(u+A_{\pi}z)-f(u+A_{p}z)\bigr{)}\,p(u,y^{\dagger}-z)\,\mathrm{d}u\,\mathrm{d}z$
$=\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}f(v)\Bigl{(}p(v-A_{\pi}z,y^{\dagger}-z)-p(v-A_{p}z,y^{\dagger}-z)\Bigr{)}\,\mathrm{d}v\,\mathrm{d}z,$ (B.28)
where we used a change of variables in the second equality. Since $\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty$ by assumption, it follows from Lemma A.3 and from the inequality $\min\{a,b\}\leqslant\sqrt{a}\sqrt{b}$ that there is a constant $C$ so that
@@
$\qquad\leqslant C\lvert A_{\pi}z-A_{p}z\rvert\cdot\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}$
×$\qquad\qquad\times\frac{1}{1+\lvert y^{\dagger}-z\rvert^{\max\{3+d_{u},4+d_{y}\}}}.$
We apply this inequality to bound for fixed $z\in^{d_{y}}$ the inner integral in (B.28). Considering only the terms that depend on $v$ gives
$\int_{{}^{d_{u}}}\lvert f(v)\rvert\max\left\{\frac{1}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}},\frac{1}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\right\}\,\mathrm{d}v$ (B.29)
$\qquad\leqslant\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v+\int_{{}^{d_{u}}}\frac{\lvert f(v)\rvert}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v$
$\qquad\leqslant\int_{{}^{d_{u}}}\frac{1+\lvert v\rvert^{2}}{1+\lvert v-A_{\pi}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v+\int_{{}^{d_{u}}}\frac{1+\lvert v\rvert^{2}}{1+\lvert v-A_{p}z\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}v$
$\qquad\leqslant\int_{{}^{d_{u}}}\frac{1+\lvert w+A_{\pi}z\rvert^{2}}{1+\lvert w\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}w+\int_{{}^{d_{u}}}\frac{1+\lvert w+A_{p}z\rvert^{2}}{1+\lvert w\rvert^{\max\{3+d_{u},4+d_{y}\}}}\,\mathrm{d}w,$
where we used that $\lvert f(v)\rvert\leqslant 1+\lvert v\rvert^{2}$, as well as a change of variable in the last line. It follows that the integral in (B.29) is bounded from above by
@@
where the inequality follows from (B.25). Finally, the resulting integral in the $z$ variable can be bounded analogously, which gives
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma B.12], where a complete proof is given. ∎
@@
In [le2009large] machinery is established to prove Monte Carlo error estimates between the finite particle ensemble Kalman filter and its mean field limit. We use such results as a component in proving Theorem 2.4 and, in so-doing, explicit dependence on moments must be tracked. In this section we give a self-contained proof of [le2009large, Theorem 5.2], following the analysis closely [^1] The result of [le2009large] holds more generally for functions $\Psi$ that are locally Lipschitz and that grow at most polynomially at infinity [le2009large, Assumption B]. However, our proof uses that $\Psi$ is globally Lipschitz, for instance in (D.12). and, in addition, tracking dependence on moments; this moment dependence may be useful in the context of future work generalizing what we do in this paper. This leads to the following error estimate stating the desired Monte Carlo error estimate.
@@
where $u_{j+1}^{(i)}$ evolve according to the iteration in (2.5). Suppose that the data $Y^{\dagger}=\{y_{j}^{\dagger}\}_{j=1}^{J}$ and the matrices $(\Sigma,\Gamma)$ satisfy Assumptions (H1) and (H2). We assume that the vector field $\Psi$ satisfies Assumption (H3) and that $h$ is a linear transformation, i.e. that Assumption (V2) is satisfied and hence also Assumption (H4). Furthermore, if the vector field $\Psi$ additionally satisfies $|\Psi|_{C^{0,1}}\leqslant\ell_{\Psi}<\infty$, then for all $\phi$ satisfying Assumption P1, there exists a constant $C=C\bigl{(}\mathcal{M}_{q}(\mu^{{\rm EK}}_{0}),J,R_{\phi},L_{\phi},\varsigma,\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{\Psi},\Sigma,\Gamma\bigr{)}$ where $q\coloneqq\max\{4^{J+1},4\varsigma,2(\varsigma+1)\}$ such that
$\left(\mathbb{E}~{}\Bigl{|}\mu^{{\rm EK},N}_{J}[\phi]-\mu^{{\rm EK}}_{J}[\phi]\Bigr{|}^{2}\right)^{\frac{1}{2}}\leqslant\frac{C}{\sqrt{N}}.$
@@
To prove the proposition, we apply the coupling argument used in [le2009large]. Using similar notation to the one in [le2009large], to the interacting $N$-particle system $\bigl{\{}u^{(i)}_{j}\bigr{\}}_{n=1}^{N}$ evolving according to the ensemble Kalman dynamics (2.5), we couple $N$ copies of the mean field dynamics $\bigl{\{}\overline{u}^{(j)}_{j}\bigr{\}}_{n=1}^{N}$ evolving according to the mean field ensemble Kalman dynamics (2.1). The mean field replicas are synchronously coupled to the interacting particle system, in the sense that they are initialized at the same initial condition and driven by the same noises; namely, the two particle systems are initialized at i.i.d. samples $u_{0}^{(i)}=\overline{u}_{0}^{(i)}\sim\mu_{0}^{{\rm EK}}$ for $i=1,\ldots,N$. For simplicity of notation, the forecast particles are denoted by the letter $v$, and we drop the hat notation from the forecast and simulated observations. Furthermore, we add a bar $\overline{\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}}$ to all the variables related to the synchronously coupled mean field particles, including the probability measures $\overline{\pi}^{{\rm EK}}_{j}$. We also define for $\pi\in\mathcal{P}(^{d_{u}\times d_{y}})$ the Kalman gain
@@
Note that the synchronously coupled particles are independent and identically distributed. With this set-up we proceed with the proof. Let $\phi$ satisfy Assumption P1. By applying the triangle inequality, we deduce that